# The Riemann Hypothesis is false

### v39

@@
where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
uniformly for $1\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
@@
We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$. For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))$ hence
@@
$=\frac{\Lambda(m)}{em^{it}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{it}}{r}+O(r/m^{2})\Big{)}$ (16)
$=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}+O(\varepsilon^{2}\log m).$ (17)
@@
For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{r}{m}(1+O(r/m))$. Thus by arguing as above, one arrives at

### v40

@@
where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. For $m\neq n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
uniformly for $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
@@
We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$. For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
@@
$=\frac{\Lambda(m)}{em^{it}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{it}}{r}+O(1/m)\Big{)}$ (16)
$=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}+O(\varepsilon\log m).$ (17)
@@
For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$. Thus by arguing as above, one arrives at